# Federated Prototype-based Contrastive Learning for Privacy-Preserving Cross-domain Recommendation

### v1

@@ Abstract
Cross-domain recommendation (CDR) aims to improve recommendation accuracy in sparse domains by transferring knowledge from data-rich domains. However, existing CDR methods often assume the availability of user-item interaction data across domains, overlooking user privacy concerns. Furthermore, these methods suffer from performance degradation in scenarios with sparse overlapping users, as they typically depend on a large number of fully shared users for effective knowledge transfer. To address these challenges, we propose a Federated Prototype-based Contrastive Learning (CL) method for Privacy-Preserving CDR, named FedPCL-CDR. This approach utilizes non-overlapping user information and prototypes to improve multi-domain performance while protecting user privacy. FedPCL-CDR comprises two modules: local domain (client) learning and global server aggregation. In the local domain, FedPCL-CDR clusters all user data to learn representative prototypes, effectively utilizing non-overlapping user information and addressing the sparse overlapping user issue. It then facilitates knowledge transfer by employing both local and global prototypes returned from the server in a CL manner. Simultaneously, the global server aggregates representative prototypes from local domains to learn both local and global prototypes. The combination of prototypes and federated learning (FL) ensures that sensitive user data remains decentralized, with only prototypes being shared across domains, thereby protecting user privacy. Extensive experiments on four CDR tasks using two real-world datasets demonstrate that FedPCL-CDR outperforms the state-of-the-art baselines.
Index Terms:
Contrastive Learning, Federated Learning, Prototype, and Cross-Domain Recommendation
I Introduction
@@
Cross-domain recommendation (CDR) has emerged as a critical strategy to address the challenge of data sparsity in recommendation systems by transferring knowledge, such as user-item interaction histories and review texts, across domains [1, 2]. Based on different recommendation scenarios, existing CDR can be divided into two categories: single-target CDR and multi-target CDR. The first genre [3, 4] aims to improve recommendation performance in the target domain by utilizing rich information from the source domain. However, it cannot enhance model performance across multiple domains simultaneously. To address this issue, multi-target CDR [5, 6] has emerged. Literature focusing on feature aggregation [3, 7] and feature disentanglement [8, 9] has been proposed to further improve recommendation performance. Feature aggregation methods typically learn representations in each domain separately and then design an aggregation function to combine these representations. On the other hand, feature disentanglement approaches concentrate on separating domain-common and domain-specific embeddings and transferring domain-common embeddings across domains. Although these works have achieved good performance, they still face two major challenges.
CH1. How to effectively protect user privacy when transferring knowledge across domains? Most existing literature [7, 10, 11] cannot solve this problem well. They usually assume that the user-item ratings or representations are directly transferred across domains. However, in the setting of privacy-preserving, these methods are unsuitable [12]. To resolve this problem, some privacy-preserving CDR methods [13, 14, 15] have gained lots of attention. For example, PriCDR [13] first introduces Differential Privacy (DP) technology to publish the source rating matrix and subsequently conduct CDR modeling. P2FCDR [14] independently learns embeddings in each domain using orthogonal functions and applies the local differential privacy (LDP) technique to protect these embeddings. However, these approaches introduce a trade-off between privacy preservation and recommendation accuracy, as the added noise can distort the underlying data patterns.
@@
CH2. How to improve recommendation performance with few overlapping users across domains? Many existing CDR methods depend on fully overlapping users as a bridge to transfer knowledge across domains [8, 7]. For example, as shown in Figure 1, John is an overlapping user with interactions in both the Movie and Book domains. If knowledge transfer is solely based on fully overlapping users, John may be interested in the comedic book “Bossypants” because he watches movies in the same genre, like “Anchorman.” However, in real-world datasets, there are very few overlapping users. For instance, the Amazon dataset has only a 5% overlapping user ratio [11]. As a result, the model performance will be degraded with such sparse overlapping users. To address this problem, we can utilize non-overlapping user information to improve recommendation performance. For instance, Lily shares similar interests with John based on their common interaction with the movie “Anchorman”. By effectively utilizing Lily’s interests, we can infer that John may like the thriller book “Gone Girl” based on Lily’s preference for the thriller movie “Get Out”, which couldn’t be realized by relying solely on fully overlapping users.
@@
To address these challenges, we propose FedPCL-CDR, a federated prototype-based contrastive learning (CL) approach for privacy-preserving CDR (PPCDR). Our method consists of two key modules, i.e., local domain learning and global server aggregation, where user-item interaction histories and review texts are stored in local domains and knowledge is transferred via prototypes within the federated learning (FL) framework, ensuring user privacy. Specifically, in the local domain, we first learn comprehensive user and item embeddings from user-item interactions and review texts. Subsequently, we utilize k-means clustering [16] to generate prototypes (cluster centroids). On one hand, these prototypes convey overlapping and non-overlapping user preferences as they are derived from the interest alignments of all entities in the domain. On the other hand, these prototypes serve as generalized representations of group-level preferences, making it more difficult for attackers to infer sensitive information about individual users. We then select representative prototypes based on overlapping users and upload them to the global server. The global server, in turn, models both local and global prototypes by aggregating these representative prototypes and transmits them back to the respective local domains. To effectively transfer knowledge, the local domain refines user embeddings by using both local and global prototypes in a CL manner. This dual-prototype approach allows for transferring knowledge at varying granularities, enabling more nuanced learning of user embeddings from different perspectives.
In summary, our proposed model makes the following contributions:
@@
  • We introduce a novel federated prototype-based CL approach for PPCDR that aims to solve the sparse overlapping user and privacy protection concerns.
  • Our method effectively leverages non-overlapping users for knowledge transfer by clustering all users, thereby enhancing model performance in scenarios with sparse overlapping users across domains.
@@
  • We utilize local and global prototypes to transfer knowledge within the FL framework, protecting user privacy.
  • Extensive experiments on four CDR tasks from two large-scale real-world datasets, Amazon and Douban, demonstrate the effectiveness of FedPCL-CDR compared to state-of-the-art baselines.
II Related Work
II-A Cross Domain Recommendation
Cross-domain recommendation (CDR) aims to address the challenge of data sparsity by transferring cross-domain knowledge [17]. The core step in CDR involves designing an effective transfer method to improve recommendation accuracy in sparse domains. With the advancements in deep learning, various transfer methods have emerged in CDR, including learning mapping functions across domains [18], feature combination [7], feature alignment [19], and transfer methods based on Graph Neural Networks (GNNs) [20, 21]. However, these approaches depend on fully overlapping users for knowledge transfer and overlook concerns about user privacy leakage. In this paper, we propose FedPCL-CDR to address these challenges.
II-B Privacy-Preserving CDR
With the increasing attention on user privacy, many scholars have begun studying PPCDR methods [13, 14, 22, 23, 12, 24, 25, 26]. PriCDR [13] and PPGenCDR [22] both use DP to protect user-item ratings during knowledge transfer. P2FCDR [14] is a privacy-preserving federated framework that learns an orthogonal mapping matrix to transform embeddings across domains and applies LDP to the transformed embeddings for user privacy protection. Meanwhile, PPCDR [12] introduces a federated graph framework that utilizes FL and LDP technologies to protect user privacy. Despite their effectiveness, these PPCDR methods must strike a balance between utility and privacy. We propose a prototype-based FL framework to address these limitations.
II-C Contrastive Learning
@@
Contrastive Learning (CL) has been widely used in computer vision [27, 28] and natural language processing [29, 30]. It is a self-supervised learning technique that aims to maximize the mutual information between two representations. To achieve this, InfoNCE [31] is proposed to learn representations by contrasting positive pairs (similar samples) against negative pairs (dissimilar samples), which discovers the semantic information shared by different views. Nowadays, CL has been applied to the CDR to improve representation learning [32, 19, 33, 11]. For instance, DCCDR [19] leverages CL to learn domain-specific and domain-invariant representations. Meanwhile, CL-DTCDR [11] utilizes CL to learn more representative user and item embeddings with user-item interaction data and side information. However, these methods directly utilize user-item ratings or representations to construct positive and negative pairs across domains, which is not feasible under privacy-preserving constraints. In this work, we construct prototype-based CL tasks to transfer cross-domain user interests, thereby protecting user privacy.
III Methodology
III-A Definitions and Notations
@@
We assume there are $M$ domains (clients) denoted as $\{D^{1},D^{2},...,D^{M}\}$ and a global server, where $D^{i}$ denotes the i-th domain. Within each domain, there exists a user set $U^{i}$ and an item set $V^{i}$. There are partial overlapping users and non-overlapping items across domains. The overlapping user set is denoted as $U^{o}$. Let $\textbf{R}^{i}\in\{0,1\}^{|U^{i}|\times|V^{i}|}$ represent the binary user-item interaction matrix, which is private and cannot be shared across domains. Figure 2 depicts the overall framework of FedPCL-CDR. We illustrate the paradigm for domain $D^{i}$, and the corresponding paradigm for other domains can be easily inferred accordingly.
@@
Figure 2: The framework of FedPCL-CDR. (1) Local Domain Learning: It focuses on mining non-overlapping user information by clustering all user data and transferring user interests across domains using local and global prototypes in a CL manner; It can be further divided into three components: (a) Graph Representation Learning: Introducing LightGCN to learn comprehensive user and item embeddings. (b) Clustering and Prototype Selection: Clustering all user embeddings to obtain prototypes and then select representative prototypes based on overlapping users. (c) Prototype-based Contrastive Learning: Facilitating knowledge transfer across domains in a contrastive manner using both local and global prototypes. (2) Global Server Aggregation: It aggregates representative prototypes uploaded from multiple domains to generate local and global prototypes.
III-B Local domain Learning Module
III-B1 Graph Representation Learning
@@
Inspired by the effectiveness of GNNs in capturing high-dimensional and complex relationships between users and items, we adopt LightGCN [34] to learn embeddings for user and item IDs as well as their review texts. We construct a graph $G^{i}$, where nodes represent users and items, and edges indicate interactions between them. By utilizing the graph convolution and propagation layers of LightGCN, we encode user and item embeddings based on $G^{i}$. Specifically, we denote ID embeddings and review text embeddings at the $l$-th layer as $\textbf{E}_{l}^{i(id)}$ and $\textbf{E}_{l}^{i(rev)}$, respectively. Initially, ID embeddings $\textbf{E}_{0}^{i(id)}$ are randomly initialized, while review text embeddings $\textbf{E}_{l}^{i(rev)}$ are learned using the document embedding model Doc2Vec [35]. Given the graph $G^{i}$, $\textbf{E}_{l}^{i(id)}$ and $\textbf{E}_{l}^{i(rev)}$ are calculated as follows:
@@
where D is a diagonal matrix and A is a adjacency matrix. After $l$ times of propagation, we can generate the final user and item ID embedding matrices $\textbf{E}_{u}^{i(id)}$ and $\textbf{E}_{v}^{i(id)}$ by concatenating multiple embedding matrices from $\textbf{E}_{0}^{i(id)}$ to $\textbf{E}_{l}^{i(id)}$. Similarly, we obtain the user and item review text embedding matrix $\textbf{E}_{u}^{i(rev)}$ and $\textbf{E}_{v}^{i(rev)}$. Finally, we concatenate the ID embeddings and review text embeddings to learn comprehensive user and item embeddings:
$\textbf{E}_{u}^{i}=f(\textbf{E}_{u}^{i(id)};\textbf{E}_{u}^{i(rev)});\textbf{E}_{v}^{i}=f(\textbf{E}_{v}^{i(id)};\textbf{E}_{v}^{i(rev)}),$ (2)
where $f$ represents the concatenation function. Here, we use the element-wise sum aggregation method.
III-B2 Clustering and Prototype Selection
In this subsection, our objective is to generate representative prototypes by clustering the user embeddings $\textbf{E}_{u}^{i}$. The clustering process incorporates data from all users within a domain, not limited to overlapping users. This not only leverages the shared knowledge among overlapping users but also explores the effective utilization of knowledge from non-overlapping users.
We begin by introducing the k-means algorithm [16], which aims to cluster user embeddings into $K$ groups. We obtain the cluster centroid set with $K$ clusters:
$T^{i}=\{\textbf{t}_{j}^{i}\}_{j=1}^{K}=Kmeans(\textbf{E}_{u}^{i}),$ (3)
Each cluster is defined by its centroid, acting as the central representation for that particular group. These centroids are regarded as prototypes, providing a comprehensive integration of information from similar users. As a result, we derive the prototype set $T^{i}=\{\textbf{t}_{1}^{i},\textbf{t}_{2}^{i},...,\textbf{t}_{K}^{i}\}$.

### v2

# FedPCL-CDR: A Federated Prototype-based Contrastive Learning Framework for Privacy-Preserving Cross-domain Recommendation

@@ Abstract
Cross-domain recommendation (CDR) aims to improve recommendation accuracy in sparse domains by transferring knowledge from data-rich domains. However, existing CDR approaches often assume that user-item interaction data across domains is publicly available, neglecting user privacy concerns. Additionally, they experience performance degradation with sparse overlapping users due to their reliance on a large number of fully shared users for knowledge transfer. To address these challenges, we propose a Federated Prototype-based Contrastive Learning (CL) framework for Privacy-Preserving CDR, called FedPCL-CDR. This approach utilizes non-overlapping user information and differential prototypes to improve model performance within a federated learning framework. FedPCL-CDR comprises two key modules: local domain (client) learning and global server aggregation. In the local domain, FedPCL-CDR first clusters all user data and utilizes local differential privacy (LDP) to learn differential prototypes, effectively utilizing non-overlapping user information and protecting user privacy. It then conducts knowledge transfer by employing both local and global prototypes returned from the server in a CL manner. Meanwhile, the global server aggregates differential prototypes sent from local domains to learn both local and global prototypes. Extensive experiments on four CDR tasks across Amazon and Douban datasets demonstrate that FedPCL-CDR surpasses SOTA baselines. We release our code at https://github.com/Lili1013/FedPCL_CDR.
keywords:
Privacy-Preserving , Contrastive Learning , Federated Learning , Prototype
journal: Nuclear Physics B
[…]
[first]organization=School of Electrical and Data Engineering, University of Technology Sydney,addressline=15, Broadway, Ultimo, city=Sydney, postcode=2000, state=NSW, country=Australia
1 Introduction
@@
Cross-domain recommendation (CDR) offers an effective solution to the data sparsity problem in recommendation systems by enabling knowledge transfer across different domains (Hu et al., 2018). Based on different recommendation scenarios, existing CDR can be divided into two categories: single-target CDR and multi-target CDR. The first genre (Zhu et al., 2019, 2022) aims to improve recommendation performance in the target domain by utilizing rich information from the source domain. However, it cannot enhance model performance across multiple domains simultaneously. To address this issue, multi-target CDR (Zhu et al., 2023a, b) has emerged, where each domain can act as either a source domain or a target domain. It leverages data from multiple domains to simultaneously enhance recommendation performance across all domains. These two kinds of methods apply some technologies, such as feature aggregation (Zhu et al., 2019, 2020) and feature disentanglement (Cao et al., 2022; Guo et al., 2023), to improve recommendation performance. Feature aggregation methods typically learn representations in each domain separately and then design an aggregation function to combine these representations. On the other hand, feature disentanglement approaches concentrate on separating domain-invariant and domain-specific representations and transferring domain-invariant representations to other domains. Despite their promising performance, these methods still face two major challenges.
CH1. How to effectively protect user privacy when transferring cross-domain knowledge? Most existing methods (Zhu et al., 2020; Xu et al., 2023; Lu et al., 2023) assume direct transfer of user-item ratings or representations across domains, making them unsuitable for privacy-preserving settings (Tian et al., 2024). To address this, privacy-preserving CDR (PPCDR) methods (Chen et al., 2022, 2023a; Liu et al., 2024) have gained lots of attention. For instance, PriCDR (Chen et al., 2022) uses Differential Privacy (DP) to release the source rating matrix for CDR modeling, while P2FCDR (Chen et al., 2023a) applies Local Differential Privacy (LDP) and Federated Learning (FL) to protect user embeddings. However, these methods provide limited privacy protection, as attackers can exploit external information to infer sensitive data. For example, attackers can exploit a public dataset to learn reference embeddings and correlate them with differentially private embeddings to infer user preferences.
@@
CH2. How to improve recommendation performance with sparse overlapping users across domains? Many existing CDR methods depend on fully overlapping users as a bridge to transfer cross-domain knowledge (Cao et al., 2022; Zhu et al., 2020). For example, as shown in Figure 1, John is an overlapping user with interactions in both the Movie and Book domains. If knowledge transfer is solely based on fully overlapping users, John may like the comedic book “Bossypants” because he watches movies in the same genre, like “Anchorman.” However, in real-world datasets, there are very few overlapping users. For instance, the Amazon dataset has only a 5% overlapping user ratio (Lu et al., 2023). As illustrated in Figure 1, non-ovelapping users like Lily and Bob have interactions in only one domain, making it impossible to directly transfer their preferences learned from other domains. As a result, the model performance will be degraded with such sparse overlap.
@@
To address these challenges, we propose FedPCL-CDR, a federated prototype-based contrastive learning (CL) framework for PPCDR. It consists of two key modules, i.e., local domain learning and global server aggregation, where user-item interaction histories and review texts are stored in local domains and knowledge is transferred via differential prototypes within the FL framework, ensuring user privacy. Specifically, in the local domain, we first utilize k-means clustering (MacQueen, 1967) to generate prototypes. On one hand, these prototypes convey overlapping and non-overlapping user preferences as they are derived from the interest alignments of all entities in the domain. Based on these prototypes, non-overlapping user preferences can be transferred across domains to improve recommendation performance. For example, in Figure 1, Lily shares similar interests with John based on their common interaction with the movie “Anchorman”. By effectively utilizing Lily’s interests, we can infer that John may like the thriller book “Gone Girl” based on Lily’s preference for the thriller movie “Get Out”, which couldn’t be realized by relying solely on fully overlapping users. On the other hand, these prototypes serve as generalized representations of group-level preferences, making it more difficult for attackers to infer sensitive information about individual users. To further enhance privacy, we employ LDP technology to learn differential prototypes, which are then uploaded to the global server. The global server, in turn, models both local and global prototypes by aggregating these differential prototypes and sends them back to the respective local domains. To effectively transfer knowledge, the local domain refines user embeddings by using both local and global prototypes in a CL manner. This dual-prototype approach allows for transferring knowledge at varying granularities, enabling more nuanced learning of user embeddings from different perspectives.
To summarize, we highlight four main contributions of this work:
  • 1.
@@
    We propose a novel federated prototype-based CL framework for PPCDR that aims to address the sparse overlapping user and privacy protection concerns.
  • 2.
    We explicitly leverage non-overlapping users for cross-domain knowledge transfer via clustering-based prototypes. This addresses the critical limitation of sparse overlapping users in real-world scenarios. In addition, we introduce a dual-prototype contrastive mechanism (local and global prototypes) to align user embeddings across domains at varying granularities, which captures finer-grained semantic relationships, enabling more robust knowledge transfer.
  • 3.
@@
    We utilize differential prototypes to transfer cross-domain knowledge within the FL framework, achieving stronger privacy protection.
  • 4.
    Extensive experiments on four CDR tasks, conducted on Amazon and Douban datasets, show that FedPCL-CDR outperforms SOTA baselines.
2 Related Work
In this section, we provide a brief overview of related research work, including cross-domain recommendation, privacy-preserving CDR, and contrastive learning.
2.1 Cross Domain Recommendation
Cross-domain recommendation (CDR) focuses on addressing the data sparsity problem by transferring knowledge across diverse domains or platforms. A core challenge in CDR is designing effective knowledge transfer methods from data-rich to data-sparse domains to improve recommendation accuracy. Recent deep learning approaches have explored various transfer techniques, including domain mapping (Man et al., 2017), feature combination (Zhu et al., 2020), feature alignment (Zhang et al., 2023), and graph-based methods (Cui et al., 2020; Zhao et al., 2019). For example, EMCDR (Man et al., 2017) introduces a MLP network to capture the non-linear mapping relationships across domains. GA-DTCDR (Zhu et al., 2020) first learns user and item embeddings through separate heterogeneous graphs within the domain and then combines these cross-domain embeddings using an attention network. DCCDR (Zhang et al., 2023) is designed to disentangle domain-invariant and domain-specific features while aligning the domain-invariant features across domains. HeroGRAPH (Cui et al., 2020) proposes a heterogeneous graph framework that unifies multiple domains into a shared graph structure, enabling cross-domain recommendation through meta-path-based neighbor aggregation and relation-aware attention mechanisms. However, these approaches assume open cross-domain data sharing, neglecting privacy concerns. In addition, they rely on fully overlapping users or items to connect different domains and transfer knowledge, leading to performance degradation with sparse overlapping entities. In this paper, we leverage non-overlapping user information and differential prototypes within the FL framework to address the sparse overlapping problem and mitigate the risk of user privacy leakage.
2.2 Privacy-preserving CDR
With the enactment of privacy protection laws and the increasing focus on user privacy, many scholars have begun studying PPCDR methods. These approaches leverage technologies, such as LDP (Chen et al., 2022, 2023b; Gao et al., 2019; Wang et al., 2024) and FL (Tian et al., 2024; Liu et al., 2023; Yan et al., 2022), to protect users’ sensitive information when transferring knowledge across domains. For example, PriCDR (Chen et al., 2022) employs Differential Privacy (DP) to release the source domain’s rating matrix. P2FCDR (Chen et al., 2023b) utilizes the LDP technique to add noise to the transformed embeddings before transfer. PPGenCDR (Liao et al., 2023) further leverages adversarial methods to generate fake ratings for transfer to the target domain. FPPDM (Liu et al., 2023) and FedCDR (Yan et al., 2022) introduce the FL framework to protect user privacy through distributed learning. $P^{2}$ (Lin et al., 2024) introduces a dual-target CDR framework that utilizes a private set intersection algorithm and prototype-based FL to protect user privacy. Among these, $P^{2}$ is the most closely related to our work. However, our method introduces critical advancements that address key limitations: (1) $P^{2}$ primarily relies on overlapping users to learn prototypes and transfer knowledge, which limits its effectiveness in sparse-overlap scenarios. In contrast, we explicitly leverage non-overlapping users by clustering all users (both overlapping and non-overlapping) to extract generalized prototypes. This allows knowledge transfer even in scenarios with extremely limited overlapping users, significantly improving robustness in real-world sparse settings; (2) $P^{2}$ transfer cross-domain knowledge through Constrained Dominant Set (CDS) propagation, which primarily optimizes intra-domain user embeddings using averaged prototypes. Our work, however, introduces prototype-based CL, aligning user embeddings with both local prototypes and global prototypes. This dual-prototype strategy captures finer-grained semantic relationships, improving transfer effectiveness.
2.3 Contrastive Learning
@@
Contrastive Learning (CL) has been widely used in computer vision (Chen et al., 2020; He et al., 2020a) and natural language processing (Gao et al., 2021; Chuang et al., 2022). It is a self-supervised learning technique that aims to maximize the mutual information between two representations. To achieve this, InfoNCE (Oord et al., 2018) is proposed to learn representations by contrasting positive pairs (similar samples) against negative pairs (dissimilar samples), which discovers the semantic information shared by different views. Nowadays, CL has been applied to the recommendation field to improve representation learning (Wu et al., 2021; Zhang et al., 2023; Chen et al., 2023c; Lu et al., 2023). For instance, DCCDR (Zhang et al., 2023) leverages CL to learn domain-specific and domain-invariant representations. Meanwhile, CL-DTCDR (Lu et al., 2023) utilizes CL to learn more representative user and item embeddings with user-item interaction data and side information. However, these methods directly utilize user-item ratings or representations to construct positive and negative pairs across domains, which is not feasible under privacy-preserving constraints. In this work, we employ differential prototypes to transfer user interests in a CL manner, thereby protecting user privacy.
3 Methodology
3.1 Definitions and Notations
@@
We assume there are $M$ domains (clients) and a global server, where $D^{i}$ denotes the i-th domain. Within each domain, there exists a user set $U^{i}$ and an item set $V^{i}$. There are partial overlapping users, denoted as $U^{o}$. Let $\textbf{R}^{i}\in\{0,1\}^{|U^{i}|\times|V^{i}|}$ represent the binary user-item interaction matrix. We illustrate the paradigm for domain $D^{i}$, and the corresponding paradigm for other domains can be easily inferred accordingly.
3.2 Overall Framework
@@
Figure 2 depicts the overall framework of FedPCL-CDR. It mainly includes two key modules: (1) Local Domain Learning Module: This module is designed to effectively utilize non-overlapping user information and transfer protected user interests across domains, which can be further divided into three components: (a) Graph Representation Learning: Introducing LightGCN to learn user and item embeddings. (b) Differential Prototype Learning: Applying unsupervised k-means clustering to user embeddings to generate prototypes and then utilizing LDP to learn differential prototypes. (c) Prototype-based Contrastive Learning: Facilitating knowledge transfer across domains in a CL manner using both local and global prototypes. (2) Global Server Aggregation Module: This module focuses on modeling local and global prototypes by aggregating differential prototypes based on overlapping users.
Figure 2: The framework of FedPCL-CDR. (1) Local Domain Learning: (a) Graph Representation Learning: Introducing LightGCN to learn comprehensive user and item embeddings. (b) Differential Prototype Learning: Clustering all user embeddings and applying LDP to obtain differential prototypes. (c) Prototype-based Contrastive Learning: Facilitating knowledge transfer across domains in a contrastive manner using both local and global prototypes. (2) Global Server Aggregation: It aggregates differential prototypes uploaded from multiple domains to generate local and global prototypes.
3.3 Local domain Learning Module
3.3.1 Graph Representation Learning
@@
Inspired by GNNs’ ability to model complex user-item relationships, we use LightGCN (He et al., 2020b) to learn embeddings for user and item IDs as well as their review texts. We construct a graph $G^{i}$, where nodes represent users and items, and edges indicate interactions between them. By utilizing the graph convolution and propagation layers of LightGCN, we encode user and item embeddings based on $G^{i}$. Specifically, we denote ID embeddings and review text embeddings at the $l$-th layer as $\textbf{E}_{l}^{i(id)}$ and $\textbf{E}_{l}^{i(rev)}$, respectively. Initially, ID embeddings $\textbf{E}_{0}^{i(id)}$ are randomly initialized, while review embeddings $\textbf{E}_{0}^{i(rev)}$ are learned using the document embedding model Doc2Vec (Le and Mikolov, 2014). The ID embeddings and review embeddings at the $l$-th layer are calculated as:
@@
where D is a diagonal matrix and A is an adjacency matrix. After $l$ propagation steps, the final user and item ID embedding matrices $\textbf{E}_{u}^{i(id)}$ and $\textbf{E}_{v}^{i(id)}$ are generated by concatenating multiple embedding matrices from $\textbf{E}_{0}^{i(id)}$ to $\textbf{E}_{l}^{i(id)}$. Similarly, we obtain the user and item review text embedding matrix $\textbf{E}_{u}^{i(rev)}$ and $\textbf{E}_{v}^{i(rev)}$. Finally, we concatenate ID and review embeddings to learn comprehensive user and item embeddings:
$\textbf{E}_{u}^{i}=f_{c}(\textbf{E}_{u}^{i(id)};\textbf{E}_{u}^{i(rev)});\textbf{E}_{v}^{i}=f_{c}(\textbf{E}_{v}^{i(id)};\textbf{E}_{v}^{i(rev)}),$ (2)
where $f_{c}$ represents the concatenation function. Here, we use the element-wise sum aggregation method.
3.3.2 Differential Prototype Learning
We begin by introducing the k-means algorithm (MacQueen, 1967), which aims to cluster user embeddings $\textbf{E}_{u}^{i}$ into $K$ groups. The cluster centroids serve as prototypes, representing the central characteristics of each user group with similar features. The selection of k-means clustering is motivated by its simplicity, efficiency, and proven effectiveness in grouping high-dimensional data. Unlike hierarchical or density-based methods, k-means scales well with large datasets and provides deterministic cluster assignments, which are crucial for generating stable prototypes. We obtain the prototype set:
$T^{i}=\{\textbf{t}_{j}^{i}\}_{j=1}^{K}=Kmeans(\textbf{E}_{u}^{i}).$ (3)
The clustering process incorporates data from all users within a domain, not limited to overlapping users. This not only leverages the shared knowledge among overlapping users but also explores the effective utilization of knowledge from non-overlapping users.